OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5

OpenStudy (compassionate):

First distribute the 3s.

OpenStudy (anonymous):

wat then?

OpenStudy (compassionate):

Then you have: −3(x + 3) = −3(x + 1) − 5 -3x - 9 = -3x - 3 - 5 Now you want to add like terms

OpenStudy (anonymous):

wouldnt it be -9=-8 to prove its a false equation?

OpenStudy (compassionate):

-3x - 9 = -3x - 3 - 5 0 - 9 = -8 Correct.

OpenStudy (anonymous):

thx